Graph each of the following equations:
x^2+y^=5
x+y=1
Use the graph to approximate their intersection. Use algebra to find the exact coordinates of intersection.
Since both variables have the same exponent. Would this not be a circle? Just curious.. Thankyou Maria

Answer by stanbon(75887)

Yes, x^2+y^2=5 is a circle with its center at (0,0) and having
radius = sqrt(5).
---------------
x+y=1 is a line
--------------
Cheers,
Stan H.